OpenStudy (anonymous):

help! 4 divided by 4 + radical 7 and 3 divided by radical 5 - 2

OpenStudy (anonymous):

\[\frac{ 4 }{ 4 + \sqrt{7} }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 4 + \sqrt{7} } (\frac{ 4 - \sqrt{7} }{ 4 - \sqrt{7} })\]

OpenStudy (anonymous):

\[\frac{ 4(4-\sqrt{7}) }{ 4^{2} + 4\sqrt{7} - 4\sqrt{7} - \sqrt{7}^{2} }\]

OpenStudy (anonymous):

\[\frac{ 16 - 4\sqrt{7} }{ 16 - 7 } = \frac{ 16 - 4\sqrt{7} }{ 9 }\]

OpenStudy (anonymous):

thanks!